A plane can fly 680 miles in the same time as it takes a car to go 280 miles. If the car travels 100 mph slower than the plane, find the speed of the plane.
Answer by fractalier(6550)

Starting by calling the rate of the car r, the rate of the plane r+100 and with rt=d, we can write
rt = 280
or t = 280/r
(r+100)t = 680
or t = 680/(r+100)
Set these equal and get
280/r = 680/(r+100)
Now cross-multiply and solve
680r = 280r + 28000
400r = 28000
r = 70 mph
r+100 = 170 mph, the rate of the plane...